3. Write an equation in slope-intercept form of the line with the given properties.
slope -7, passes through (7,4)
Please help me determine

Answer by Shana-D77(132)

When you're given a slope and a point, it's best to use the point-slope form of the equation. "The" equation means the linear (straight line) equation. The point slope form is:


y = y1 = m(x - x1).
((side note: this is actually the slope formula, just rearranged.))


So,


slope (m) = -7
point (7, 4)


y - y1 = m(x - x1)
(the y and the x always stay as y and x in this form)


Plug in what your teacher gave you. y1 = 4 and x1 = 7 (from (7,4)), and slope (m) = -7


y - 4 = -7(x - 7)
y - 4 = -7x + 49 (distributed the -7)
y = -7x + 53 (added 4 to both sides)